In the problem, a square with a circumference of 5 is inside of a square. The answer to the problem is 1/4 the area not covered by the circle. What is the answer and how did you get it.: In the problem, a square with a circumference of 5 is inside of a square. The answer to the problem is 1/4 the area not covered by the circle. What is the answer and how did you get it.
Answer by nabla(409)

C=pi*diameter
thus pi*d=5
and d=5/pi
d^2 will be the area of the square.
d^2=25/(pi)^2
Now, pi*r^2=Area of circle
and r=1/2d=(1/2)(5/pi)=(5/(2pi))
Thus, A= pi*(25/(4pi^2))=25/(4pi)
The left over area will be 5/(pi)^2-25/(4pi).